# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v2

@@ 5.1 Numerical Demonstration
To test our algorithms for calculating natural gradient directions, we train a neural network with about $7.5\times 10^{5}$ parameters to learn a rapidly oscillating function on the unit interval. We take three snapshots from the training process and form the least-squares problem (32) for each. The methods TA-mSGD, TA-RBK-U, and TA-ReBlocK-U are then compared on these problems with results in Fig. 5. The computations are performed on an A100 GPU to simulate a deep learning setting, and progress is measured with wall-clock time on the horizontal axis. TA-ReBlocK-U performs best in every case, achieving an accuracy over an order of magnitude better the other methods in both the second and third snapshots.
In Fig. 6, we investigate the effect of the parameter $\lambda$ on the performance of ReBlocK. We observe that choosing $\lambda=1e-1$ already provides a significant speed-up relative to mSGD, and the speed-up is further improved by reducing $\lambda$ down to $31e-$ 3. However, choosing $\lambda=1e-4$ results in unstable behavior similar to the RBK algorithm. These results suggest that the performance of ReBlocK is optimized by choosing $\lambda$ as small as possible without introducing significant instabilities. Next, in Fig. 7 we explore the effect of the batch size $k$ on the performance of the algorithms. We observe that ReBlocK performs well across all batch sizes, although RBK is competitive for the largest batch size of $k=200$. Finally, in Fig. 8, we report the number of iterations per second for mSGD, RBK, and ReBlocK for the experiments of this section. We find that the ReBlocK iterations are over twenty-five times faster than the RBK iterations and only about $6\%$ slower than the mSGD iterations, which is a significant reason that TA-ReBlocK-U outperforms TA-RBK-U in practice. Further details on these experiments can be found in F.2.
Figure 6: Performance of TA-ReBlocK-U on the problem from the top right panel of Fig. 5 with various values of $\lambda$.
@@
Our results, though preliminary due to their synthetic nature, suggest that ReBlocK is a promising method for calculating natural gradient directions. This provides justification for the Kaczmarz-inspired SPRING algorithm [19] and suggests a family of related methods to be explored by future works. Open questions include how many ReBlock iterations to run between each update of $\theta$, whether to incorporate averaging, and how to tune $\lambda$.
@@ 6 Conclusions
In this work, we have explored the problem of solving large-scale linear least-squares problems without preprocessing the data matrix $A$. Our results suggest that ReBlocK is a more effective algorithm than either RBK or mSGD for inconsistent problems that exhibit rapid singular value decay. More broadly, our work highlights the value of incorporating regularization as a path towards broadening the applicability of block Kaczmarz methods, and our analysis suggests that a Monte Carlo perspective can be useful in elucidating the behavior of randomized block row-access methods in general. Finally, our work provides motivation and suggests new directions for least-squares based natural gradient optimizers.
@@ Acknowledgements
G.G. acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (J.H., L.L.). L.L. is a Simons Investigator in Mathematics. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley. We thank Ethan Epperly, Robert Webber, Ruizhe Zhang, and anonymous reviewers for their thoughtful discussions and feedback.
@@ Appendix A Helpful Lemmas
In this appendix we prove five lemmas which will help us analyze both RBK and ReBlocK. Lemma A.1 shows how the expectation of the iterates evolves. Lemma A.3 shows that the iterates of the algorithms always stay within $\operatorname{range}(A^{\top})$, and Lemma A.5 provides a contraction property of applying $I-\overline{P}$ to the certain types of vectors. Lemma A.7 shows how to derive bounds on the expected square error of tail-averaged iterates under relevant assumptions. Finally, Lemma A.9 bounds the attained residual based on the condition number of $\overline{W}$.
@@
First expand $x$ in the basis of eigenvectors of the symmetric matrix $\overline{P}$, then note that every eigenvector that has a nonzero coefficient in the expansion has its eigenvalue in the interval $[\alpha,1]$.
@@
(39) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}B+V$
for all $t$ and some constants $B,V$. Then the tail averages $\overline{x}_{T}$ of the iterates, with burn-in time $T_{b}$, satisfy
(40) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{1}{\alpha(T-T_{b})}V.$
@@
We follow closely the Proof of Theorems 1.2 and 1.3 of [18]. Decompose the expected mean square error as
@@
(42) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}.$
Apply another course bound
(43) $\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}\leq 2\sum_{r=T_{b}+1}^{T}\sum_{s=0}^{\infty}(1-\alpha)^{s}=\frac{T-T_{b}}{\alpha}$
@@
(44) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{\alpha(T-T_{b})}.$
Lemma A.9 (Residual bound).
@@
(45) $\left\lVert r^{(\rho)}\right\rVert^{2}\leq\kappa(\overline{W})\left\lVert r\right\rVert^{2}.$
@@
Use operator norms and the optimality of $x^{(\rho)}$ for the weighted least-squares problem:
@@ Appendix B Proofs of RBK Convergence Theorems
In this section we provide the proofs of Theorems 3.1, 3.2, and 3.3. We first prove the following lemma, which can be viewed as a more general version of Theorem 1.2 of [31]:
@@
Iterating from $s=0$ to $s=t-1$ and utilizing $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$ yields the desired result.
@@
Furthermore, Lemma B.1 holds and provides the conditions for Lemma A.7 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$. The application of Lemma A.7 then directly implies the convergence result for the tail-averaged RBK-U algorithm.
@@
In the case of Gaussian data, suppose we are given $m$ data points $\{a_{i},b_{i}\}_{i=1}^{m}$. Then the dataset $A,b$ can be viewed as a finite sample of the underlying statistical linear regression problem
(51) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$
Furthermore, the solution $x^{*}$ is the corresponding (minimal-norm) empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ [25, 44].
Now, the RBK-U algorithm is applied by uniformly selecting $k$ samples per iteration from the finite data $A,b$. Nonetheless, in the limiting case where $m\to\infty$, the sampling distribution of $k$ rows from the finite dataset $\{a_{i},b_{i}\}_{i=1}^{m}$ converges to the distribution obtained by sampling $k$ items independently from $\mathcal{N}(0,Q)$. Thus, to obtain results about the RBK-U algorithm in the limit $m\rightarrow\infty$, we can focus on the case in which the $k$ data points $\{a_{1},b_{1}\},\dots,\{a_{k},b_{k}\}$ are directly sampled from $\mathcal{N}(0,Q).$
@@
Denote by $Q=VV^{\top}$ the Cholesky decomposition of $Q$. Let $V=\begin{pmatrix}L\\ d^{\top}\end{pmatrix}$, where :::: and $d^{\top}$ are the top $n$ rows and the last row of $V$, respectively. Recalling that $Q_{n}$ represents the top left $n\times n$ block of $Q$, it follows that $Q_{n}=LL^{\top}$ is the Cholesky decomposition of $Q_{n}$, as previously defined.
@@
In each iteration of RBK-U, the $k$ data points $[a_{1}^{\top}\;b_{1}],\dots,[a_{k}^{\top}\;b_{k}]$ from $\mathcal{N}(0,Q)$ can be equivalently redistributed as $z_{1}^{\top}[L^{\top}\;d],\dots,z_{k}^{\top}[L^{\top}\;d]$ for $z_{i}\sim\mathcal{N}(0,I_{n+1})$. Collecting the random vectors $z_{1},\ldots,z_{k}$ into the columns of a single matrix $Z_{t}$, the RBK-U update can be written as
(52) $x_{t+1}=x_{t}+(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}d-Z_{t}^{\top}L^{\top}x_{t}).$
@@
Now, note that $\mathbb{E}\left[a_{i}a_{i}^{\top}\right]=LL^{\top},\mathbb{E}\left[b_{i}a_{i}\right]=Ld$ and $\mathbb{E}\left[b_{i}^{2}\right]=d^{\top}d$. Plugging these identities into the definition of $y$, we find
(53) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right]=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\|L^{\top}x-d\|^{2}.$
@@
We can thus define an “underlying” residual vector by $\tilde{r}=L^{\top}y-d$. Additionally defining $P_{t}=(Z_{t}^{\top}L^{\top})^{+}Z_{t}^{\top}L^{\top}$, the update (52) can be reformulated as
(54) $x_{t+1}-y=(I-P_{t})(x_{t}-y)+(Z_{t}^{\top}L^{\top})^{+}Z_{t}^{\top}\tilde{r}.$
Noting from the definition that $\tilde{r}$ is orthogonal to the rows of $L$, we now show $\mathbb{E}\left[(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}\tilde{r})\right]=0$. In fact, decompose $Z_{t}=Z_{t}^{1}+Z_{t}^{2}$ where $Z_{t}^{1}\in\operatorname{range}(L^{\top})$ and $Z_{t}^{2}\perp\operatorname{range}(L^{\top})$. It follows that $Z_{t}^{1}$ and $Z_{t}^{2}$ are independent, zero-mean Gaussian vectors and
$\mathbb{E}\left[(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}\tilde{r})\right]=\mathbb{E}\left[((Z_{t}^{1})^{\top}L^{\top})^{+}(Z_{t}^{2})^{\top}\tilde{r}\right]=\mathbb{E}\left[((Z_{t}^{1})^{\top}L^{\top})^{+}\right]\mathbb{E}\left[(Z_{t}^{2})^{\top}\tilde{r}\right]=0.$
@@
We note that a similar independence lemma has also been established in Lemma 3.14 of [37]. Furthermore, taking the expectation on both sides of (54), we have
@@
where $\overline{P}:=\mathbb{E}\left[P_{t}\right]$. Consequently, in the infinite limit $\mathbb{E}\left[x_{t}\right]$ converges to $y$ using the same logic as the proof of Theorem 3.1, which proves that
(55) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@
Now, consider the bound on the value of $\alpha$. Note that the entries of $Z_{t}\in\mathbb{R}^{d\times m}$ are independently drawn from the standard Gaussian distribution. This property enables us to leverage existing results from [14] concerning the spectrum of the matrix $\overline{P}$. Specifically, $P_{t}=(Z_{t}L^{\top})^{+}Z_{t}L^{\top}$, which has a similar formula to the matrix $P$ considered in Equation (6) of [14]. Then, applying Theorem 3.1 of [14] gives the result stated in Theorem 3.2.
@@
It remains to show the bound on the variance. Note that the variance term $\left\lVert A_{S_{t}}^{\top}r_{S_{t}}\right\rVert^{2}$ in Theorem 3.1 takes the form
(56) $\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}.$
We can thus bound
(57) $\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}=\left\lVert(Z_{1}^{\top}L^{\top})^{+}Z_{2}^{\top}\tilde{r}\right\rVert^{2}\leq\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{-2}\left\lVert Z_{2}^{\top}\tilde{r}\right\rVert^{2}.$
@@
We can bound the expectation of the smallest singular value of $Z_{1}^{\top}L^{\top}$ using a similar technique to the proof of Lemma 22 in [14]. To start, let $L^{\top}=W\Sigma Y^{\top}$ be the SVD of $L^{\top}$. Then
(58) $\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}=\sigma_{\rm min}(Z_{1}^{\top}L^{\top}LZ_{1})=\sigma_{\rm min}(Z_{1}^{\top}W\Sigma^{2}W^{\top}Z_{1}).$
@@
Letting $W_{2k}$ denote the first $2k$ columns of $W$, and noting that $\Sigma\succeq{\rm diag}(\sigma_{2k},\dots,\sigma_{2k},0,\dots,0)$, it holds
$\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}\geq\sigma_{2k}^{2}\cdot\sigma_{\rm\min}(Z_{1}^{\top}W_{2k}W_{2k}^{\top}Z_{1})=\sigma_{\min}(Z_{1}^{\top}W_{2k}).$

### v3

@@ 6.1 Numerical Demonstration
Figure 9: Per iteration plots for the same three snapshots from Figure 8, with TA-RBK-U run for much longer than the other methods to produce a comparable number of iterations. The algorithms are measured in terms of their progress towards reducing the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$ for the least-squares problem (34), which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. The burn-in period is set based on seconds rather than iterations and so is not indicated in this plot.
To test our algorithms for calculating natural gradient directions, we train a neural network with about $7.5\times 10^{5}$ parameters to learn a rapidly oscillating function on the unit interval. We take three snapshots from the training process and form the least-squares problem (34) for each. The methods TA-mSGD, TA-RBK-U, and TA-ReBlocK-U are then compared on these problems with results in Fig. 8. The computations are performed on an A100 GPU to simulate a deep learning setting and in double precision to enable each algorithm’s iterations to be computed accurately, and progress is measured with wall-clock time on the horizontal axis. TA-ReBlocK-U performs best in every case. Interestingly, RBK-U fails catastrophically on the first snapshot and also exhibits instabilities in the last snapshot, providing further evidence that the issues highlighted by the no-go Equation 18 are not merely of theoretical interest. We also provide per iteration convergence plots in Figure 9. In the first snapshot the catastrophic failure of TA-RBK-U persists while in the second and third snapshots, we find that TA-RBK-U is more competitive on a per iteration basis, but is still outperformed by TA-ReBlocK-U.
In Fig. 10, we investigate the effect of the parameter $\lambda$ on the performance of ReBlocK. We observe that choosing $\lambda=1$ already provides an advantage relative to mSGD, and the advantage is further improved by reducing $\lambda$ down to 1e31e-3 or 1e61e-6. However, when choosing $\lambda=1e-9$ the variance becomes so large as to significantly hinder the performance of the algorithm. These results suggest that the performance of ReBlocK is optimized by choosing $\lambda$ as small as possible without introducing significant instabilities. Next, in Fig. 11 we explore the effect of the batch size $k$ on the performance of the algorithms. We observe that ReBlocK performs best across all batch sizes. Finally, in Fig. 12, we report the number of iterations per second for mSGD, RBK, and ReBlocK for the experiments of this section. We find that the ReBlocK iterations are over twenty-five times faster than the RBK iterations and only about $10\%$ slower than the mSGD iterations, which provides a significant boost to TA-ReBlocK-U even beyond its improved stability. Further details on these experiments can be found in Appendix G.
Figure 10: Performance of TA-ReBlocK-U on the problem from the top left panel of Fig. 8 with various values of $\lambda$.
@@
Our results, though preliminary due to their synthetic nature, suggest that ReBlocK is a promising method for calculating natural gradient directions. This provides justification for the Kaczmarz-inspired SPRING algorithm [goldshlager2024kaczmarz] and suggests a family of related methods to be explored by future works. Open questions include how many ReBlocK iterations to run between each update of $\theta$, whether to incorporate averaging, and how to tune $\lambda$.
@@ 7 Conclusions
In this work, we have explored the problem of solving large-scale linear least-squares problems without preprocessing the data matrix $A$. Our results suggest that ReBlocK is a more effective algorithm than either RBK or mSGD for inconsistent problems that exhibit rapid singular value decay. More broadly, our work highlights the value of incorporating regularization as a path towards broadening the applicability of block Kaczmarz methods, and our analysis suggests that a Monte Carlo perspective can be useful in elucidating the behavior of randomized block row-access methods in general. Finally, our work provides motivation and suggests new directions for both natural gradient optimizers and time stepping methods in scientific machine learning.
@@ Acknowledgements
We thank Ethan Epperly, Robert Webber, Marius Zeinhofer, Ruizhe Zhang, and the anonymous reviewers for their thoughtful discussions and feedback.
@@ Appendix A Helpful Lemmas
In this appendix we prove five lemmas which will help us analyze both RBK and ReBlocK. Lemma A.1 shows how the expectation of the iterates evolves. Lemma A.3 shows that the iterates of the algorithms always stay within $\operatorname{range}(A^{\top})$, and Lemma A.5 provides a contraction property of applying $I-\overline{P}$ to certain types of vectors. Lemma A.7 shows how to derive bounds on the expected square error of tail-averaged iterates under relevant assumptions. Finally, Lemma A.9 bounds the weighted residual and the bias based on the condition number of $\overline{W}$.
@@
First expand $x$ in the basis of eigenvectors of the symmetric matrix $\overline{P}$, then note that every eigenvector that has a nonzero coefficient in the expansion has its eigenvalue in the interval $\alpha,1]$.
@@
(41) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}B+\tilde{V}$
for all $t$ and some constants $B,\tilde{V}$, where we use $\tilde{V}$ to distinguish from the variance term $V$ in Theorems 3.1 and 5.1. Then the tail averages $\overline{x}_{T}$ of the iterates, with burn-in time $T_{b}$, satisfy
(42) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{2}{\alpha(T-T_{b})}\tilde{V}$.
@@
We follow closely the Proof of Theorems 1.2 and 1.3 of [epperly2024randomizedkaczmarztailaveraging]. Decompose the expected mean square error as
@@
(44) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{\tilde{V}}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}.$
Apply another coarse bound
(45) $\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}\leq 2\sum_{r=T_{b}+1}^{T}\sum_{s=0}^{\infty}(1-\alpha)^{s}=\frac{2(T-T_{b})}{\alpha}$
@@
(46) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{2}{\alpha(T-T_{b})}\tilde{V}$.
Lemma A.9 (Residual and bias bounds).
@@
(47) $\left\lVert r^{(\rho)}\right\rVert^{2}$ $\leq\kappa(\overline{W})\cdot\left\lVert r\right\rVert^{2},$
$\left\lVert x^{(\rho)}-x^{*}\right\rVert$ $\leq\sqrt{\kappa(\overline{W})-1}\cdot\left\lVert A^{+}\right\rVert\cdot\left\lVert r\right\rVert.$
@@
To bound the weighted residual, use operator norms and the optimality of $x^{(\rho)}$ for the weighted least-squares problem:
@@
Next, note that the weighted residual admits an orthogonal decomposition $r^{(\rho)}=A(x^{*}-x^{(\rho)})+r$. It follows that
(48) $\left\lVert A(x^{(\rho)}-x^{*})\right\rVert^{2}=\left\lVert r^{(\rho)}\right\rVert^{2}-\left\lVert r\right\rVert^{2}\leq(\kappa(\overline{W})-1)\cdot\left\lVert r\right\rVert^{2}.$
Taking square-roots and using $\left\lVert x^{(\rho)}-x^{*}\right\rVert\leq\left\lVert A^{+}\right\rVert\cdot\left\lVert A(x^{(\rho)}-x^{*})\right\rVert$ yields the bias bound
(49) $\left\lVert x^{(\rho)}-x^{*}\right\rVert\leq\sqrt{\kappa(\overline{W})-1}\cdot\left\lVert A^{+}\right\rVert\cdot\left\lVert r\right\rVert.$
@@ Appendix B Proofs of RBK Convergence Theorems
In this section we provide the proofs of Theorems 3.1, 4.1, and 4.2. We first prove the following lemma, which can be viewed as a more general version of Theorem 1.2 of [needell2014paved]:
@@
Iterating from $s=0$ to s=t1s=t-1 and utilizing $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$ yields the desired result.
@@
Furthermore, Lemma B.1 holds and provides the conditions for Lemma A.7 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $\tilde{V}=\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$. The application of Lemma A.7 then directly implies the convergence result for the tail-averaged RBK-U algorithm.
Now, we proceed with upper bounds on condition number, the weighted residual, the bias, and the variance. For the condition number, first note that
(55) $\left\lVert A_{S}A_{S}^{\top}\right\rVert=\left\lVert A_{S}^{\top}A_{S}\right\rVert=\left\lVert\sum_{i\in S}a_{i}a_{i}^{\top}\right\rVert\leq\sum_{i\in S}\left\lVert a_{i}a_{i}^{\top}\right\rVert\leq k\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2}.$
Since we have assumed the rows of $A$ are in general position, this implies $\|(A_{S}A_{S}^{\top})^{+}\|\geq 1/(k\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2})$. Since $\overline{W}=\mathbb{E}_{S}\left[I_{S}^{\top}(A_{S}A_{S}^{\top})^{+}I_{S}\right]$ follows that
$\frac{1}{k\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2}}\cdot\mathbb{E}[I_{S}^{\top}I_{S}]\preceq\overline{W}\preceq\max_{S}\|A_{S}^{+}\|^{2}\cdot\mathbb{E}[I_{S}^{\top}I_{S}].$
When $S$ is sampled uniformly, $\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}I_{S}\right]=\frac{k}{m}I$ and so this implies
$\kappa(\overline{W})\leq\ k\cdot\Big(\max_{i}\|a_{i}\|^{2}\Big)\cdot\max_{S}\|A_{S}^{+}\|^{2}.$
The bounds on the weighted residual and the bias follow directly from Lemma A.9. Finally, regarding the variance, it follows from the definition of $V$ that
(56) $V$ $=\mathbb{E}\,_{S}\|A_{S}^{+}r_{S}^{(\rho)}\|^{2}\leq\max_{S}\|A_{S}^{+}\|^{2}\ \cdot\,\mathbb{E}_{S}\|r_{S}^{(\rho)}\|^{2}$
$=\max_{S}\|A_{S}^{+}\|^{2}\cdot\frac{k\|r^{(\rho)}\|^{2}}{m}.$
Applying the weighted residual bound from Lemma A.9 the desired result.
@@
In the case of Gaussian data, we consider the statistical linear regression problem
(57) $x^{*}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a^{\top}\;b]\sim\mathcal{N}(0,Q)}\left[(a^{\top}x-b)^{2}\right].$
@@
The RBK-U algorithm is applied by generating $k$ independent samples per iteration from the data distribution $\mathcal{N}(0,Q)$. Denote by $Q=LL^{\top}$ the Cholesky decomposition of $Q$. Let $L=\begin{bmatrix}L_{a}\\ \ell_{b}^{\top}\end{bmatrix}$ so $L_{a}$ represents the first $n$ rows of $L$ and $\ell_{b}^{\top}$ represents the last row of $L$. Due to the lower triangular nature of $L$ it follows that $L_{a}=\begin{bmatrix}L_{n}&0\end{bmatrix}$ where $L_{n}$ is the $n\times n$ Cholesky factor of $Q_{n}$ (which we previously defined to be the top left $n\times n$ block of $Q$).
@@
In each iteration of RBK-U, the $k$ data points $a_{1}^{\top}\;b_{1}],\dots,[a_{k}^{\top}\;b_{k}]$ from $\mathcal{N}(0,Q)$ can be equivalently redistributed as $z_{1}^{\top}[L_{a}^{\top}\;\ell_{b}],\dots,z_{k}^{\top}[L_{a}^{\top}\;\ell_{b}]$ for $z_{i}\sim\mathcal{N}(0,I_{n+1})$. Collecting the random vectors $z_{1},\ldots,z_{k}$ into the columns of a single matrix $Z_{t}$, the RBK-U update can be written as
(58) $x_{t+1}=x_{t}+(Z_{t}^{\top}L_{a}^{\top})^{+}(Z_{t}^{\top}\ell_{b}-Z_{t}^{\top}L_{a}^{\top}x_{t}).$
@@
Now, note that $\mathbb{E}\left[aa^{\top}\right]=L_{a}L_{a}^{\top},\mathbb{E}\left[ba\right]=L_{a}\ell_{b}$ and $\mathbb{E}\left[b^{2}\right]=\ell_{b}^{\top}\ell_{b}$. Plugging these identities into the definition of $x^{*}$, we find
(59) $x^{*}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a^{\top}\;b]\sim\mathcal{N}(0,Q)}\left[(a^{\top}x-b)^{2}\right]=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\|L_{a}^{\top}x-\ell_{b}\|^{2}.$
@@
We can thus define an “underlying” residual vector by $\tilde{r}=\ell_{b}-L_{a}^{\top}x^{*}$. Additionally defining $P_{t}=(Z_{t}^{\top}L_{a}^{\top})^{+}Z_{t}^{\top}L_{a}^{\top}$, the update (58) can be reformulated as
(60) $x_{t+1}-x^{*}=(I-P_{t})(x_{t}-x^{*})+(Z_{t}^{\top}L_{a}^{\top})^{+}Z_{t}^{\top}\tilde{r}$.
Now, decompose $Z_{t}^{\top}=[Z_{t}^{1}\;Z_{t}^{2}]$ where $Z_{t}^{1}$ consists of the first $n$ columns and $Z_{t}^{2}$ of the last column. Note that $Z_{t}^{1},Z_{t}^{2}$ are independent mean-zero Gaussian matrices and note also that that since $L_{a}=[L_{n}\;0]$ it holds $Z_{t}^{\top}L_{a}^{\top}=Z_{t}^{1}L_{n}^{\top}$. Furthermore since $L_{n}$ is full rank, $\operatorname{range}(L_{a}^{\top})={\rm span}(e_{1},\ldots,e_{n})$ and thus $\tilde{r}\parallel e_{n+1}$ (here $e_{i}$ is the $i$th standard basis vector in $\mathbb{R}^{n+1}$). It follows that $Z_{t}^{\top}\tilde{r}=Z_{t}^{2}\tilde{r}_{n+1}$ and so
$\mathbb{E}\left[(Z_{t}^{\top}L_{a}^{\top})^{+}(Z_{t}^{\top}\tilde{r})\right]=\mathbb{E}\left[(Z_{t}^{1}L_{n}^{\top})^{+}Z_{t}^{2}\tilde{r}_{n+1}\right]=\mathbb{E}\left[(Z_{t}^{1}L_{n}^{\top})^{+}\right]\mathbb{E}\left[Z_{t}^{2}\tilde{r}_{n+1}\right]=0$
@@
since $\mathbb{E}\left[Z_{t}^{2}\right]=0$. We note that a similar independence lemma has also been established in Lemma 3.14 of [rebrova2021block]. Furthermore, taking the expectation on both sides of (60), we have
@@
where $\overline{P}:=\mathbb{E}\left[P_{t}\right]$. Consequently, $\mathbb{E}\left[x_{t}\right]$ converges to $x^{*}$ using the same logic as the proof of Theorem 3.1, which proves that $x^{(\rho)}=x^{*}$.
@@
Now, consider the bound on the value of $\alpha$. Note that the entries of $Z_{t}\in\mathbb{R}^{(n+1)\times k}$ are independently drawn from the standard Gaussian distribution. This property enables us to leverage existing results from [derezinski2024sharp] concerning the spectrum of the matrix $\overline{P}$. Specifically, $P_{t}=(Z_{t}L_{a}^{\top})^{+}Z_{t}L_{a}^{\top}$, which has a similar formula to the matrix $P$ considered in Equation (6) of [derezinski2024sharp]. Then, noting that $L_{a}=[L_{n}\;0]$ has the same singular value spectrum as $L_{n}$, applying Theorem 3.1 of [derezinski2024sharp] gives the result stated in Theorem 4.1.
@@
It remains to show the bound on the variance. Note that the variance term $\left\lVert A_{S_{t}}^{\top}r_{S_{t}}\right\rVert^{2}$ in Theorem 3.1 here takes the form
(61) $\left\lVert(Z_{t}^{\top}L_{a}^{\top})^{+}Z_{t}^{\top}\tilde{r}\right\rVert^{2}.$
We bound
(62) $\left\lVert(Z_{t}^{\top}L_{a}^{\top})^{+}Z_{t}^{\top}\tilde{r}\right\rVert^{2}=\left\lVert(Z_{t}^{1}L_{n}^{\top})^{+}Z_{t}^{2}\tilde{r}_{n+1}\right\rVert^{2}\leq\sigma_{\rm min}^{-2}(Z_{t}^{1}L_{n}^{\top})\left\lVert Z_{t}^{2}\tilde{r}_{n+1}\right\rVert^{2}.$
@@
We can bound the expectation of the smallest singular value of $Z_{t}^{1}L_{n}^{\top}$ using a similar technique to the proof of Lemma 22 in [derezinski2024sharp]. To start, let $L_{n}^{\top}=W\Sigma Y^{\top}$ be the SVD of $L_{n}^{\top}$. Then
(63) $\sigma_{\rm min}^{2}(Z_{t}^{1}L_{n}^{\top})=\sigma_{\rm min}(Z_{t}^{1}L_{n}^{\top}L_{n}(Z_{t}^{1})^{\top})=\sigma_{\rm min}(Z_{t}^{1}W\Sigma^{2}W^{\top}(Z_{t}^{1})^{\top}).$
@@
Let $W_{2k}$ denote the first $2k$ columns of $W$ and note that $\Sigma\succeq{\rm diag}(\sigma_{2k},\dots,\sigma_{2k},0,\dots,0)$ where we recall that $\sigma_{i}$ is the $i$th singular value of $L_{n}$. It thus holds
$\sigma_{\rm min}^{2}(Z_{t}^{1}L_{n}^{\top})\geq\sigma_{2k}^{2}\cdot\sigma_{\rm\min}(Z_{t}^{1}W_{2k}W_{2k}^{\top}(Z_{t}^{1})^{\top})=\sigma_{\min}^{2}(Z_{t}^{1}W_{2k}).$